OpenStudy (rulerlol113):

a -4 = what i'm confused here.

OpenStudy (rulerlol113):

\[a ^{-4}\]

OpenStudy (photon336):

That means a to the negative fourth power. We can rewrite this as 1/a^4

OpenStudy (madhu.mukherjee.946):

@Rulerlol113 what is the problem?

OpenStudy (adrimit):

\[a^{-k}=(a^{-1})^{k}=\left(\frac{1}{a}\right)^{k}=\frac{1}{a^k}\]

OpenStudy (rulerlol113):

thanks